An inclined plane is 100 m long and at an angle of 20 degrees to the





horizontal. The AMA of the slope is two. Calculate:





a) the effort required to push a 72 underline C N block up the slope b) the ideal mechanical advantage c) the efficiency of the slope.

1
Expert's answer
2022-03-22T11:00:37-0400

Explanations & Calculation


  • Mechanical advantage of an incline is

"\qquad\qquad
\begin{aligned}
\small MA&=\small \frac{mg}{F}\\
\end{aligned}"

  • Using this,

"\qquad\qquad
\begin{aligned}
\small F_{actual}&=\small \frac{72\,N}{2}=36\,N\\
\\
\small IMA&=\small \frac{mg}{mg\sin\theta}=\frac{1}{\sin\theta}=\frac{1}{\sin20}\\
&=\small 2.9(>AMA)
\end{aligned}"